Creator: Gisli Ottarsson <gisli@liapunov.eecs.umich.edu>

Description: Calvin and Hobbes

\begin{matrix}k_{c_{3}} \\ k_{c_{3}} \\ k_{c_{1}} \\ k_{c_{2}} \end{matrix}

Creator: leecheng@liapunov.eecs.umich.edu

Description: Dripping faucet model.

Dripping Faucet  ^{M m  g Mathematical Model for a Dripping Faucet}

Creator: Christian Schytt <pierre@diku.dk>

Description: Primal and dual.

c_{ij}  p_{i} − p_{j} Dual cost ; of arc (i, j) q_{ij}(p_{i} − p_{j})  Slope: − l_{ij} Slope: − u_{ij} ^{l_{ij}  u_{ij}  f_{ij} Primal ; cost ; of arc (i, j) a  Slope: c_{ij}}

Creator: Gisli Ottarsson <gisli@liapunov.eecs.umich.edu>

(r_{o}, θ_{o})  x_{o}  u(r_{o}, θ_{o})  w(x_{o})

Creator: tvz Description: Another example of \pspolygon. The coordinates

where determined using \psgrid, after making the table.

Result is true for values in shaded region:

X11 X12 X13 X14 X15 X16
X21 X22 X23 X24 X25 X26
X31 X32 X33 X34 X35 X36
X41 X42 X43 X44 X45 X46

Creator: tvz Description: Nodes.

Returns to Scale Production  Managing Supervision  Information Processing

Creator: tvz Description: Another example of nodes.

0  1  2  3_{start}_{a}_{b}_{b}_{b}_{b}_{a}_{a}_{a}

Creator: tvz Description: Nonsense examples of nodes.

The tempestuous Maggie and her alcoholic husband Brick …

Much Ado about Nothing.

and another another another another another book

Creator: tvz Description: More node nonsense.

GOAT

GOAT

Doom

Zoom Polar Coor Polar Coor H M r l Doom Creator: tvz Description:

\begin{aligned}U \\ & X ×_{Z} Y & X \\ & Y & Z \end{aligned} _{y}_{x}_{q}_{p}_{f}_{g}

Creator: tvz Description: Several ways to connect nodes to themselves

•  •  •

1  2  3

1  2  3

Creator: tvz Description:

A connection  Unto another

Creator: tvz Description: